OpenStudy (anonymous):

Solve the following equations: 1) x^2-47=242 2) x^3=216 3) x^6=729

OpenStudy (campbell_st):

for the 1st question add 47 to both sides of the equation, then take the square root. there are 2 answers the positive and negative, 2. take the cube root of both sides of the equation, the answer is positive 3. take the 6th root... and there will 2 answers, positive and negative

OpenStudy (anonymous):

1. x^2 = 289 x= +- 17 2. x^3 = 216 x= 6 3. x^6 = 729 x= +- 3

OpenStudy (anonymous):

thank you !!!!!!!!!! <3 Campbell: i have very short time on my test left and i really suck on mathematics i need to read the book after this !!! lol thank u guys!!!! <3